OpenStudy (anonymous):

Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x) = e^-5x [0, 1] so I did (e^-5(1))-(e^5(0))/1-0 that gave me (e^-5)-1 but when I plug that in it won't work, so I figured that since the answer ends up being -.9933 and that the interval is [1,0] it would be DNE, but that isn't right either. help?

myininaya (myininaya):

\[f'(c)=\frac{f(b)-f(a)}{b-a}\]

myininaya (myininaya):

\[-5e^{-5c}=\frac{e^{-5(1)}-e^{-5(0)}}{1-0}\]

myininaya (myininaya):

\[-5e^{-5c}=\frac{e^{-5}-1}{1}\]

myininaya (myininaya):

\[-5e^{-5c}=e^{-5}-1\]

myininaya (myininaya):

solve for c

OpenStudy (anonymous):

mmm could you explain why you set it equal to -5e^-5c?

myininaya (myininaya):

its the formula i wrote above \[e^{-5c}=\frac{e^{-5}-1}{-5}\]

myininaya (myininaya):

\[f'(x)=(-5)e^{-5x} =>f'(c)=-5e^{-5c}\]

OpenStudy (anonymous):

ahh okay thank you :)

myininaya (myininaya):

\[\ln(e^{-5x})=\ln(\frac{e^{-5}-1}{-5}) => -5x=\ln(\frac{e^{-5}-1}{-5})\]

myininaya (myininaya):

i meant that to be c not x (but whatever) \[c=\frac{-1}{5} \cdot \ln(\frac{e^{-5}-1}{-5})\] and this c is in btw 0 and 1